OpenStudy (anonymous):

-1/6(x-18)+1/3(x+3)=x-11

OpenStudy (anonymous):

First distribute the 6 in the first term -1/6x-108 Are you with me so far?

OpenStudy (anonymous):

kinda this one is alot harder

OpenStudy (anonymous):

Yes it is. Don't let the -1 on the numerator phase you out. Just focus on the denominator, which is currently 6x-108 because 6(x) - 6(18) = 6x-108

OpenStudy (anonymous):

It's not hard... cmenendez, i can explain an easy way to do it

OpenStudy (anonymous):

But you can stay here with kirby, i'm sure he will be able to help you out :)

OpenStudy (anonymous):

Then you can look at the second term, which is 1/3(x+3) = 1/3x + 3(3) =1/(3x+9) So far we have -1/(6x-108) + 1/(3x+9) = x-11 It is not simplified to a certain extent. Do you follow?

OpenStudy (anonymous):

thanks i will.

OpenStudy (anonymous):

I mean now XD it is now simplified

OpenStudy (anonymous):

Then you can make the terms have the same denominators. The goal will be (6x-108)(3x+9)